OpenStudy (anonymous):

7/8 p-4=10 What is P

OpenStudy (anonymous):

When solving two-step equations, use inverse operations to undo addition or subtraction first then multiplication or division

OpenStudy (anonymous):

so undo -4 by adding 4 to both sides

OpenStudy (anonymous):

ok now what

OpenStudy (anonymous):

7/8 p = 14

OpenStudy (anonymous):

you want to isolate y by getting one as the coefficient so one option is to multiply 7/8 byt the reciprocal--both sides

OpenStudy (anonymous):

7/8 * 8/7 = 1 so 1p or p = 14(8/7)

OpenStudy (anonymous):

p = 14/1 * 8/7

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

p=16

OpenStudy (anonymous):

p=16

OpenStudy (anonymous):

plug p=16 into original equation to check 7/8 (16) -4=10

OpenStudy (anonymous):

14-4=10 10=10 YES!

OpenStudy (anonymous):

TIS RITE

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have more

OpenStudy (anonymous):

i will in a little bit

OpenStudy (anonymous):

ok...medalplease : )

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8+3N/12=13 what is N

OpenStudy (anonymous):

N=20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

g/-5+3=-13

OpenStudy (anonymous):

find g

OpenStudy (anonymous):

sorry was on fb

OpenStudy (anonymous):

+ 3 first

OpenStudy (anonymous):

no way me to

OpenStudy (anonymous):

do you want to learn the steps

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

g/-5 = -16 (subtract 3 to undo + 3)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

coefficient of g is -1/5 so what is the repriocal(flip numerator and denominator)

OpenStudy (anonymous):

-5/1 so multiply -16 by -5 (-5/1) g = 80